# CLUE-Mark: Watermarking Diffusion Models using CLWE

### v3

@@
Abstract
As AI-generated images become widespread, reliable watermarking is essential for content verification, copyright enforcement, and combating disinformation. Existing techniques rely on heuristic approaches and lack formal guarantees of undetectability, making them vulnerable to steganographic attacks that can expose or erase the watermark. Additionally, these techniques often degrade output quality by introducing perceptible changes, which is not only undesirable but an important barrier to adoption in practice.
@@ Abstract
In this work, we introduce CLUE-Mark, the first provably undetectable watermarking scheme for diffusion models. CLUE-Mark requires no changes to the model being watermarked, is computationally efficient, and because it is provably undetectable is guaranteed to have no impact on model output quality. Our approach leverages the Continuous Learning With Errors (CLWE) problem — a cryptographically hard lattice problem — to embed watermarks in the latent noise vectors used by diffusion models. By proving undetectability via reduction from a cryptographically hard problem we ensure not only that the watermark is imperceptible to human observers or adhoc heuristics, but to any efficient detector that does not have the secret key. CLUE-Mark allows multiple keys to be embedded, enabling traceability of images to specific users without altering model parameters. Empirical evaluations on state-of-the-art diffusion models confirm that CLUE-Mark achieves high recoverability, preserves image quality, and is robust to minor perturbations such JPEG compression and brightness adjustments. Uniquely, CLUE-Mark cannot be detected nor removed by recent steganographic attacks.
1 Introduction
In an era where AI-generated images are becoming increasingly common, identifying whether an image originates from a model is critical for combating disinformation and settling copyright claims. Watermarking, wherein a signal is inserted into an image such that it can later be recovered, has emerged as an essential tool to determine the provenance of images. Current approaches to watermarking AI-generated images remain ad-hoc. For example, Tree Ring [1] was conjectured to be undetectable and robust, but no proof as such was given. While an important step forward, Tree Ring watermarks can be both detected and removed by standard steganalysis [2], i.e. averaging the difference between watermarked and non-watermarked images.
This is more than a security concern; it shows that the watermark affects the output of the model such that the output images are both quantitatively and qualitatively different from the original. After spending considerable effort to train and optimise large diffusion models, providers and users are understandably reluctant to then degrade the quality of the images they generate in order to watermark them.
@@ 1 Introduction
Thus for both security and quality reasons, it is important that the watermark be provably undetectable, by which we mean that a watermarked image be indistinguishable from an unwatermarked image. A rigorous reduction from a cryptographically hard problem is needed, otherwise it may turn out that a watermark that was conjectured to be undetectable is in fact detectable by some other means, or may affect downstream applications. By proving undetectability, we ensure that the watermark does not compromise the quality of the generated images and remains stealthy to detectors.
In this work we will show how cryptographic techniques can be used to inject a watermark into diffusion model generated images such the watermark is provably indistinguishable from the no watermark case. Previous work on watermarking AI-generated images has shown that by injecting signals into the noise inputs of diffusion models, it is possible to recover watermarks even after significant modifications to the image by inverting the model and analyzing the noise for the watermark signal [1, 3]. While these techniques demonstrated robustness to perturbations arising in benign applications, they were not provably undetectable and more recent work has shown that they can be detected and removed by steganalysis [2].
Undetectability from CLWE.
@@
Our approach is to modify the noise input to a diffusion model using a distribution based on the Continuous Learning With Errors (CLWE) problem [4]. The intuitive idea is that this distribution produces samples that appear normal in all directions except for a secret direction in which it has a signature pattern that is different from normal. By carefully constructing the watermark, we can prove that any process that can reliably detect the watermark can be used to distinguish a CLWE distribution from a normal distribution. This problem is known to be cryptographically hard [4], and thus we can be confident that CLUE-Mark is undetectable for any downstream process, not just those we have tested, ensuring image quality and undetectability. Moreover, because CLUE-Mark only modifies the input noise, it does not modify the model in any way and does not require retraining. Meanwhile, if the secret direction is known then the watermark can be efficiently recovered by inverting the model and checking for the signature.
A secondary advantage to this approach is that it supports multiple keys, and several keys can be used in a single watermark. For example, a provider may have a key for all images generated by a model and a key for each user, allowing them to quickly check if an image came from the model before checking which user generated it.
In an ideal world, the perfect scheme would be such that watermarked images can always be verified to contain the watermark (completeness) and unmarked images can always be verified to not contain the watermark (soundness), without degrading image quality, and despite perturbations to the image that do not degrade image quality. Unfortunately, we prove in Section 4 that such perfect watermarking is impossible theoretically. As such, we must choose which properties are most important for our application. In this work we aim first for undetectability that can be rigorously proven, and then secondarily for a watermark that is as robust as possible.
Empirical Results.
While CLUE-Mark is theoretically undetectable, the question becomes whether it is usable in practice. Image sampling and inversion of diffusion models adds considerable noise, and thus the watermark may not be recoverable. If so, can the watermark be recovered after applying common perturbations? To that end, we evaluate CLUE-Mark on commonly used diffusion models and datasets to verify undetectability, recoverability, image quality, and robustness. We show that CLUE-Mark is indeed undetectable by common means, produces images with the same quality as the original, and is recoverable with the secret key. We also show that it is robust to small perturbations, such as JPEG compression and brightness.
@@
Our contribution is a novel and practical watermarking scheme, CLUE-Mark, for images generated by modern diffusion AI models. It is provably undetectable, in the sense that detectability is as hard a distinguishing the CLWE distribution from a Gaussian—a standard cryptographic hardness assumption in lattice-based cryptography.
2 Motivation and Problem Setup
Figure 1: Sampling images from latent diffusion models
Diffusion models, such as Stable Diffusion [5, 6], have quickly emerged as the go-to AI models for generating very high quality and customizable images. They have been adopted across diverse fields such as digital art and content creation, e-commerce, and aiding research in natural sciences (see survey [7]). The process to generate an image from a diffusion model is called “sampling”, in which the model is given a text embedding and a noise vector (in the latent space of the model, hence this is often called the “latent vector” or “latent noise”), and the model iteratively removes the noise to produce an image described by the text embedding, as illustrated in Figure 1. We refer to [6, 5] for understanding the details of diffusion models, and describe only the parts that are necessary for our work.
@@ 2 Motivation and Problem Setup
Malicious actors can (and do) misuse diffusion models to create realistic fake images, like deepfakes, which pose serious risks. They use these images to blackmail [8], execute scams [9], spread disinformation [10], and manipulating individuals with potentially harmful financial, social, and political consequences. Watermarking is a proposed measure to detect AI-generated content. Model developers, such as OpenAI and Google, have pledged to watermark the content they generate with active monitoring and government oversight [11, 12, 13].
@@
Watermarking an image refers to implanting a signal within it such that the signal can later be recovered. If an image can be confirmed to contain the watermark signal, this offers proof of the provenance of the image. In a secret key scheme, only those with the key can verify an image’s origin. Consider a company that hosts a closed-source model. If the company watermarks the images generated by the model using a secret key, they can later verify whether an image was generated by the model by checking for the watermark using the key. They can also give the key to another entity to do verification, for example a social media platform that wants to verify the authenticity of images uploaded by users. The company may also assign a key to each user in order to trace which user was responsible for generating a particular image. Watermarks may be applied to images after generation (“post-process”) or during model execution (“in-process”). Diffusion models offer another option: watermarking the noise vector $\mathbf{z}$ used to generate the image (“pre-process”), as introduced by Tree Rings [1]. In this work we focus on secret-key pre-process watermarking. See Figure 2 for an illustration of the use cases.
Figure 2: Watermarking Use Cases.
Problem.
@@
Despite the potential benefits of watermarking AI generated images, adoption has been limited [14]. One reason for this is image quality. Large models that generate high-quality images are expensive to train and serve, and both companies and users are sensitive to any degradation in image quality. Further, changes to image quality may also affect the performance of downstream tasks. For example, if the generated images are used to train another model or as inputs to a larger process, any changes in the images may affect the performance of the system as a whole.
@@
Existing watermarking techniques for diffusion models have a noticeable impact on the images generated. Post-processing techniques, e.g. embedding a signal into the output pixel data and even AI-based techniques, have visible quality losses [1]. As we evaluate in this work, state-of-the-art techniques such as Tree Ring and Gaussian Shading, despite aiming for not being perceptible, produce images with a significant quality differences from their unwatermarked equivalents (see Section 6.4). This is because of ad-hoc design choices and the lack of provable guarantees. When evaluated over a large set of images, it can be seen that the watermarked images are biased (less diverse statistically) than the unwatermarked images. Further, recent attacks have shown that the watermarks can be detected, removed, and forged using standard steganographic techniques [2].
2.1 Undetectable Watermarking
@@
[…]
@@
In order to be useful, we require that a watermarking scheme be at least complete and sound, for which we use Definitions 2.3 and 2.4.
@@
[…]
@@
[…]
@@
Before we can define undetectability, we must define formally what we mean by “distinguishing” and how we measure success. For this we use Definition 2.5.
@@
We want to ensure that our watermarking scheme is undetectable, that is, that the output of the watermarking process is indistinguishable from non-watermarked images generated by the model. We formally define this in Definition 2.6.
@@
[…]
3 The CLUE-Mark Approach
@@
Our approach to watermarking images is based on the idea of modifying the latent noise vector of a diffusion model (i.e. pre-process watermarking). Recall that a diffusion model takes as input a latent vector that is expected to be from the standard gaussian distribution. In our approach, we instead use a latent vector that is drawn from the hCLWE distribution (which will be defined in Section 3.1). This vector is indistinguishable from the standard gaussian distribution to all PPT-computable functions without the secret key, but can be efficiently distinguished when the secret key is known. We call using hCLWE samples to generate an image the watermark injection process. To test whether an image is watermarked, a system with access to the secret key uses the watermark recovery process. During recovery, we invert the diffusion model to get an estimate of the initial latent vector from the watermarked image. We then use the secret key to determine if the latent vector is from the hCLWE distribution, and if so, then it is very likely that the image is watermarked using that key.
@@
In this section we describe the high-level approach of watermarking, including the necessary background on CLWE, and the watermark injection and recovery processes. We will prove undetectability of the construction in Section 4.
3.1 Background on CLWE

### v4

@@
Abstract.
The proliferation of AI-generated images has led to a new problem known as “AI Slop”: low-effort, mass-produced AI-generated images that are distributed as-is but are difficult to distinguish from authentic images. Watermarking such content from popular diffusion models provides a promising solution: if social media platforms and end users can easily identify images generated by the most common AI providers, then a very large proportion of the AI Slop problem can be detected by end users. However, prior watermarking techniques are heuristic and lack formal guarantees of undetectability. Techniques that were claimed to not affect image quality were later shown to in fact degrade quality, hampering adoption as providers and users are averse to any reduction in output quality.
@@ Abstract.
In this work, we introduce CLUE-Mark, a provably undetectable watermaking scheme for diffusion models. CLUE-Mark requires no changes to the model being used, is computationally efficient, and guaranteed to have no impact on model output quality. Our approach leverages the Continuous Learning With Errors (CLWE) problem — a cryptographically hard lattice problem — to embed hidden messages in the latent noise vectors used by diffusion models. We reduce the security of CLWE against an adversary given limited samples to the security of CLUE-Mark. Assuming that CLWE is secure in the limited sample regime, image watermarks generated by CLUE-Mark are undetectable by any efficient adversary that does not have the secret key, making them imperceptible to human observers and undetectable by ad-hoc heuristics as well. CLUE-Mark focuses primarily on undetectability while maintaining sufficient robustness for common non-adversarial transformations such as JPEG compression. Empirical evaluations on state-of-the-art diffusion models confirm that CLUE-Mark achieves high message recovery, preserves image quality, and is robust to minor perturbations such as JPEG compression and brightness adjustments. A proof-of-concept implementation of CLUE-Mark and source code to reproduce our evaluations is available at: https://github.com/kisp-nus/cluemark
1. Introduction
The internet has recently witnessed an unprecedented surge of AI-generated content, an effect that was first popularly referred to as “AI Slop” in mid-2024 (Hern and Milmo, [n. d.]; Hoffman, [n. d.]). This phenomenon has now entered popular discourse with coverage from diverse media outlets including The Guardian (Hern and Milmo, [n. d.]; Malik, 2025; Mahdawi, [n. d.]), The New York Times (Hoffman, [n. d.]), New York Magazine Intelligencer (Read, [n. d.]), and John Oliver’s Last Week Tonight (Oliver, 2025). As pointed out by many commentators, the problem with such content is not the generation of content itself, but that for many users it is difficult to distinguish authentic and synthetic content, leading to problems of misinformation and user trust.
Such machine-generated content is created in large volumes and shared through social media with little to no modification. Many users feel either frustrated by the inability to distinguish it from authetic content or overwhelmed by the sheer volume of it, or both, with many blaming social media platforms for doing too little to identify AI generated content (Hern and Milmo, [n. d.]; Malik, 2025). Thus, the problem of identifying AI-generated images is both timely and important.
One proposed method to identifying AI-generated images is watermarking, in which the image is “marked” in such a way that it can be distinguished from an unmarked image. In this manner, should publicly available AI image generation tools embed watermarks in their output, as has been proposed by several legal projects (ai_, 2023; of Public Affairs, 2024; Diane and Krystal, 2024; Madiega, 2023), it would be possible to identify AI-generated content automatically either by social media platforms or by end-user software. As the name and physical analogy suggests, watermarking is typically designed to be robust to benign transformations, but is also easily detectable by the user, producing at least some visible artifacts or other degradation in image quality. This presents a major barrier to adoption among both providers and users of AI image generation tools, as they are (understandably) reluctant to degrade the quality of the images generated by their models after spending considerable effort to train them.
Prior approaches to watermarking AI-generated images were ad-hoc. For example, Tree Ring (Wen et al., 2023) was conjectured to be undetectable and robust, but no proof as such was given. While an important step forward, Tree Ring watermarks can be both detected and removed by standard steganalysis (Yang et al., 2024a), i.e. averaging the difference between watermarked and non-watermarked images. Similarly, Gaussian Shading (Yang et al., 2024b) was proved to be undetectable, but only under the assumption that the nonce used to generate the watermark is not reused; otherwise the watermark is again detectable by steganalysis. The one-time (per image) nonce assumption is hard to satisfy practically. This is more than a security concern; it shows that the watermark affects the output of the model such that the output images are both quantitatively and qualitatively different from the original (a finding that we verify in Section 5.5).
@@ 1. Introduction
Thus for both security and quality reasons, it is important that the watermark be provably undetectable, by which we mean that a watermarked image must be indistinguishable from an unwatermarked image111 Embedding a signal in an image such that it cannot be detected but is not robust to perturbations is technically considered steganography rather than watermarking, but but the latter term remains popular.. A rigorous reduction from a cryptographically hard problem is needed, otherwise it may turn out that a watermark that was conjectured to be undetectable is in fact detectable by some other means, or may affect downstream applications. By proving undetectability, we ensure that the watermark does not compromise the quality of the generated images and remains stealthy to detectors. In an ideal world, such an undetectable watermark (properly: steganography) would also be robust to benign transformations. Unfortunately, it has been proved that a perfect watermark cannot exist (see Appendix A and (Zhang et al., 2024)), and thus we must choose between robustness and undetectability.
In this work we present CLUE-Mark, an undetectable watermark (i.e. steganography) scheme for diffusion models to address the problem of AI Slop. Given the challenges listed above and the impossibility of perfect watermarking, we focus on using cryptographic techniques to inject hidden information into diffusion model generated images such that any efficient algorithm that can distinguish marked and unmarked images can be used to break fundamental cryptographic assumptions — something assumed to be impossible — while verifiers that have access to the secret key can recover the hidden information with high probability, despite minor purturbations such as JPEG compression. This allows labelling of AI-generated content at scale, precise the problem posed by AI Slop. Even if sophisticated users employing more effort or advanced techniques can remove the hidden watermark, this prevents low-effort users from flooding social media networks with unmarked AI-generated content, drastically improving the scale of the problem. By proving undetectability, we ensure that our method does not compromise the quality of the generated images.
Undetectability from CLWE
@@
Our approach modifies the noise input to a diffusion model using a distribution based on the Continuous Learning With Errors (CLWE) problem (Bruna et al., 2021). Intuitively, this distribution is gaussian (i.e. normal) in all directions except for a secret direction in which it has a signature pattern. For our CLUE-Mark construction, we can prove that any process that can reliably detect this hidden pattern can be used to distinguish a CLWE distribution from a normal distribution. Assuming that this problem is cryptographically hard, we can guarantee that our approach is undetectable by any efficient algorithm without the secret key, ensuring perfect image quality and compatibility with all content creation tools. Because our technique only modifies the input noise, it does not require any model modifications or retraining.
Empirical Results
While CLUE-Mark aims for undetectability, we must also demonstrate its practical utility. Can our steganographic signal be reliably recovered after image generation and common non-adversarial transformations like JPEG compression? To answer this, we evaluate CLUE-Mark on commonly used diffusion models and datasets to verify undetectability, recoverability, image quality, and robustness to non-adversarial transformations. We show that CLUE-Mark is undetectable by standard means, produces images with identical quality to those without steganographic content, and the hidden message can be reliably recovered after JPEG compression and minor brightness adjustments. A proof-of-concept implementation of CLUE-Mark and source code to reproduce our evaluations is available at: https://github.com/kisp-nus/cluemark
Comparison to PRC-based Watermarks
In work contemporary to ours, Gunn et al (Gunn et al., 2025) propose using Pseudo-Random Codes (PRC) to bias the latent input to a diffusion model for watermarking. While PRC offer a promising new primitive, our work uses CLWE—a complementary primitive to encode secret information in the continuous $\mathbb{R}^{d}$ domain directly—which is more natural fit for machine learning systems. PRC-based watermarks use parameters that are not known to be theoretically secure, as stated in their work, and is more suitable for applications where robustness to adversarial perturbation is more important than undetectability. We discuss in more detail how our work differs from PRC watermarks in Section 6.
@@
Our contribution is a novel and practical watermarking scheme, CLUE-Mark, for images generated by modern diffusion AI models. It aims to be undetectable, in the sense that distinguishing marked and unmarked images is as hard as distinguishing the CLWE distribution from a Gaussian. This assumption is stricter than the standard cryptographic hardness of CLWE; it considers the practically-motivated regime where the adversary has limited samples. Along the way we develop a novel hCLWE sampling algorithm (see Section 4.3) that is more practical than the previous rejection-sampling technique proposed in (Bruna et al., 2021), which may be of independent interest. These tools enables detection of low-effort AI-generated content (“AI Slop”) without compromising image quality or compatibility with existing tools and platforms.
2. Motivation and Problem Setup
Figure 1. Sampling images from latent diffusion models
Diffusion models, such as Stable Diffusion (Podell et al., 2023; Rombach et al., 2022), have quickly emerged as the go-to AI models for generating high quality and customizable images. They have been adopted by content creators across diverse fields, from digital art to professional illustration for e-commerce and advertising (see survey (Yang et al., 2023)). The process to generate an image from a diffusion model is called “sampling,” in which the model is given a text embedding and a noise vector (in the latent space of the model, hence this is often called the “latent vector” or “latent noise”), and the model iteratively removes the noise to produce an image described by the text embedding, as illustrated in Figure 1. We refer to (Rombach et al., 2022; Podell et al., 2023) for understanding the details of diffusion models, and describe only the parts that are necessary for our work.
@@ 2. Motivation and Problem Setup
Watermarking is a proposed measure to detect AI-generated content, which can be used to fight against AI Slop. Watermarking an image refers to implanting a signal within it such that the signal can later be recovered, providing proof of the provenance of the image. In a secret key scheme, only those with the key can verify an image’s origin, which is technically closer to steganography than watermarking, the latter usually prioritizes robustness to perturbations over undetectability. Consider a company that hosts a closed-source model. If the company watermarks the images generated by the model using a secret key, they can later verify whether an image was generated by the model by checking for the watermark using the key. They can also give the key to another entity to do verification, for example a social media platform that wants to verify the provenance of images uploaded by users. The company may also assign a key to each user in order to trace which user was responsible for generating a particular image. Notice that the verifier, whether the model provider, social media platform, or another party, is given the model keys but not any per-image data, such as text prompts, nonces, or noise vectors. Watermarks may be applied to images after generation (“post-process”) or during model execution (“in-process”). Diffusion models offer another option: watermarking the noise vector $\mathbf{z}$ used to generate the image (“pre-process”), as introduced by Tree Rings (Wen et al., 2023). In this work we focus on secret-key pre-process watermarking. See Figure 2 for an illustration of the use cases.
Figure 2. Watermarking use cases and our threat model, with a per-model (and per-creator) keys. Note that the verifier is given model keys, but no image-specific data, such as text prompts, noise vectors, or per-image nonces or keys.
Previous approaches
@@
Despite the potential benefits of watermarking AI generated images, adoption has been limited (Seetharaman and Barnum, 2023). One reason for this is image quality. Large models that generate high-quality images are expensive to train and serve, and both companies and users are sensitive to any degradation in image quality. Further, changes to image quality may also affect the performance of downstream tasks. For example, if the generated images are used to train another model or as inputs to a larger process, any changes in the images may affect the performance of the system as a whole.
@@
Existing watermarking techniques for diffusion models have a noticeable impact on the images generated. Post-processing techniques, e.g. embedding a signal into the output pixel data and even AI-based techniques, have visible quality losses (Wen et al., 2023). As we evaluate in this work, state-of-the-art techniques such as Tree Ring and Gaussian Shading, despite aiming for not being perceptible, produce images with a significant quality differences from their unwatermarked equivalents (see Section 5.4). This is because of ad-hoc design choices and the lack of provable guarantees. When evaluated over a large set of images, it can be seen that the watermarked images are biased (less diverse statistically) than the unwatermarked images. Further, recent attacks have shown that the watermarks can be detected, removed, and forged using standard steganographic techniques (Yang et al., 2024a).
Problem
The question thus becomes: can we watermark diffusion model generated images such that a verifier as in Figure 2 can distinguish between watermarked and unwatermarked images, while proving that the watermark does not affect the quality of the images? We formally define this problem in the next section.
2.1. Undetectable Watermarking
@@
A watermarking function is defined as the tuple of functions $(\text{Setup},\text{Mark}$, $\text{Extract})$:
@@
For a watermarking scheme to be useful, it needs to satisfy both completeness and soundness criteria, for which we use Definitions 3 and 4 below. Notice that the definitions are worst-case over all prompts. Our approach currently focuses on embedding a single bit of information, though this can be extended to multiple bits through phase-shifting techniques or using multiple keys.
@@
A watermarking scheme $(\text{Setup},\text{Mark}$, $\text{Extract})$ is $\delta$-complete if for all $\pi\in\mathcal{T}^{*}$:
@@
Where the probability is taken over all random coins used in Setup, Mark, and Extract.
@@
A watermarking scheme $(\text{Setup},\text{Mark}$, $\text{Extract})$ is $\delta$-sound if for all $\pi\in\mathcal{T}^{*}$:
@@
Where the probablity is taken over $\mathbf{z}$ and all random coins used in Setup.
Before we can define undetectability, we must define formally what we mean by “distinguishing” and how we measure success. For this we use Definition 5.
@@
We want to ensure that our watermarking scheme is undetectable, that is, that the output of the watermarking process is indistinguishable from non-watermarked images generated by the model. For both users and providers, this guarantees that watermarked outputs have identitical quality to unmarked outputs, and remains compatible with all existing tools and platforms. We formally define this in Definition 6.
@@
Let $\mathcal{M}_{\theta}$ be an oracle that on input $\pi\in\mathcal{T}^{*}$ samples $\mathbf{z}\leftarrow\mathcal{N}(\mathbf{0},\mathbf{I})$ and returns $\text{Sample}_{\theta}(\pi;\mathbf{z})$. Let $\mathcal{S}_{\theta,k}$ be an oracle that on input $\pi\in\mathcal{T}^{*}$ returns $\text{Mark}_{\theta}(\pi;k)$ for a fixed key $k\leftarrow\text{Setup}_{\theta}(1^{\lambda})$. A watermarking scheme $(\text{Setup},\text{Mark}$, $\text{Extract})$ is undetectable if for all probabilistic polynomial-time PPT distinguishers $D$, $Adv_{D,\mathcal{M}_{\theta},\mathcal{S}_{\theta,k}}\leq negl(\lambda)$.
While robustness against adversarial manipulations is a secondary concern in our AI Slop prevention use case, we still require the embedded information to survive common non-adversarial transformations that occur in normal usage, such as JPEG compression when images are shared on social media platforms (see Figure 2). Our approach prioritizes absolute undetectability first, while maintaining sufficient robustness for practical usage.
3. The CLUE-Mark Approach
@@
Our approach to steganography is based on the idea of modifying the latent noise vector of a diffusion model (i.e. pre-process steganography). Recall that a diffusion model takes as input a latent vector that is expected to be from the standard gaussian distribution. In our approach, we instead use a latent vector that is drawn from the hCLWE distribution (which will be defined in Section 3.1). This vector is indistinguishable from the standard gaussian distribution to all PPT-computable functions without the secret key, but can be efficiently distinguished when the secret key is known. We call using hCLWE samples to generate an image the signal injection process. To test whether an image contains a signal, a system with access to the secret key uses the signal recovery process. During recovery, we invert the diffusion model to get an estimate of the initial latent vector from the image. We then use the secret key to determine if the latent vector is from the hCLWE distribution, and if so, then it is very likely that the image contains a signal using that key.
@@
In this section we describe the high-level approach of steganography, including the necessary background on CLWE, and the signal injection and recovery processes. We will prove undetectability of the construction in Section 4.
3.1. Background on CLWE